ROTATIONAL ANALYSIS OF A-E AND E-E VIBRONIC BANDS OF CF_{3}S

\underline{MIN-CHIEH YANG}, DAVID E. POWERS, CHRISTOPHER C. CARTER, JAMES M. WILLIAMSON, and TERRY A. MILLER, The Ohio State University, Department of Chemistry, The Laser Spectroscopy Facility, 120 West 18th Ave., Columbus, OH 43210.

Rotationally resolved spectra of jet-cooled CF_{3}S have been obtained. The rotational analysis of the origin band shows that it is an a-e type vibronic transition. This is also consistent with the Ã ^{2}A_{1} - X̃ ^{2}E electronic transition, expected for molecules of this type. This indicates that CF_{3}S has nominal C_{3v} symmetry in both its ground and first excited electronic state.

In addition to the origin band, several transitions to excited vibrational levels of the Ã state have been observed. Some of these transitions are vibronically a-e while others appear to be e-e, presumably induced by Jahn-Teller interaction in the X̃ state. The analysis of an e-e band proves to be a challenge. Its appearance is quite different from that of the a-e bands. The rotational structure of the a-e band is very congested since the entire band spans only 4 cm^{−1}. However, the e-e band spans more than 7 cm^{−1} with a concomitant decrease in line density. Most of the observed rotational lines in the e-e band belong to Q branches. The P and R branches do exist but the intensities are much weaker than the corresponding Q branches. The details of analysis will be presented.